Steel firms in the Bay Area produce steel which has a demand equal to PMB=100-Q, where Q is measured in tons. The firms have a private marginal cost of production equal to PMC=10+Q/2. This industrial activity releases numerous pollutants, some of which are deposited onto nearby land (like heavy metals), and the marginal external damage inflicted upon neighbors’ land is MD = Q/4.

What tax per ton of steel would make the market equilibrium equal to the efficient equilibrium?
Answer #1

n2 = q msc= PMC + mo - =10+ą tapo lo ta - 10 At social equilibrium, PMB = msc. 10o-Q = 10 + 3Q 90= 70 Msc = B. PMC Q = 51.43